### v1

# PatchTrack: A Comprehensive Analysis of ChatGPT’s Influence on Pull Request Outcomes

Abstract
The rapid adoption of large language models (LLMs) like ChatGPT in software development has introduced new ways for developers to interact with AI, particularly in pull request workflows. While prior research has examined AI-generated code quality, there is limited understanding of how ChatGPT is utilized in real-world pull request decision-making and how its suggestions influence patch integration and rejection. To explore these aspects, we analyze self-admitted ChatGPT usage (SACU), where developers explicitly disclose their reliance on ChatGPT within pull request discussions. Our study examines 338 pull requests (285 merged, 53 closed) across 255 GitHub repositories, containing 645 ChatGPT-generated code snippets and 3,486 patches. We introduce PatchTrack, a classification tool that determines whether ChatGPT-generated patches were applied (PA, 115 cases), not applied (PN, 64 cases), or not suggested (NE, 106 cases). Our findings reveal that full adoption of ChatGPT-generated code is rare, developers frequently modify or selectively integrate AI-generated patches to align with project constraints, with a median integration rate of 25%. Through qualitative analysis, we identify key factors influencing patch integration and pull request rejection, including scope misalignment, maintainability concerns, redundant solutions, and procedural barriers such as incomplete documentation or administrative policies. By providing empirical insights into ChatGPT’s role in pull request workflows, this study informs developers, maintainers, and educators on the evolving use of generative AI in collaborative software development. It also lays the groundwork for future research on optimizing AI-assisted development, improving transparency in AI adoption, and enhancing patch integration workflows.
keywords:
[…]
AI-assisted workflows are increasingly shaping software development, with tools such as ChatGPT and GitHub Copilot that improve automation, developer productivity, and collaboration [1, 2, 3, 4, 5, 6, 7]. Prior studies have explored their role in code generation, debugging, and documentation [8, 9, 10, 11], yet their impact on collaborative workflows, particularly in pull request decision-making, remains less understood. Understanding how developers integrate, modify, or reject AI-generated code in pull requests is essential to improve AI-assisted collaboration.
Pull requests are central to modern software engineering, enabling collaborative contributions and rigorous code review processes [12, 13, 14]. Despite their importance, PR workflows remain effort-intensive, requiring developers to evaluate code quality, resolve integration conflicts, and ensure alignment with project goals [6, 7]. The long-standing vision of the Programmer’s Apprentice [15] envisioned AI as a reasoning collaborator, offering suggestions, help with debugging and design insights, while leaving key decisions to developers. Recent advances in generative AI have brought this vision closer to reality, with tools now increasingly integrated into everyday development workflows. However, the effectiveness of such tools in practical pull request scenarios - especially to influence integration, revision, or rejection decisions - remains largely unclear [16, 17, 18, 19].
Although prior research has explored predictive models and recommendation systems for pull request decision making [20, 21, 22], AI-assisted software engineering research has focused primarily on code quality, debugging, and documentation improvements[23, 24, 25, 26]. Studies evaluating GitHub Copilot and ChatGPT highlight their productivity benefits[27, 28, 29, 30, 31, 32] but offer limited insights into collaborative pull request workflows. Furthermore, while ChatGPT-generated code has been studied in the context of integration rates and modification trends [33, 34], the decision-making processes behind patch acceptance, modification, and rejection remain largely unexplored.
This study addresses these gaps by analyzing ChatGPT conversations shared in merged and closed pull requests. We classify AI-assisted pull request outcomes and examine the factors that influence developer decisions to accept, modify, or reject ChatGPT-generated patches. Unlike previous work that focused on AI-generated code retention [33] or security concerns [34], we investigated the broader role of ChatGPT in pull request decision-making, including cases where pull requests are closed despite AI involvement. Our study builds on research by Champa et al. [18] and Chouchen et al. [35], situating ChatGPT’s role within real-world pull request workflows to assess its impact on collaboration, decision-making, and code review efficiency.
To examine ChatGPT’s influence on pull request workflows, we analyze both merged and closed pull requests where developers shared ChatGPT-generated content, including both code patches and conceptual guidance, leveraging the platform’s link-sharing feature introduced in May 2023 [36]. Unlike studies that primarily assess AI-generated code quality or developer modifications [33, 34], our study investigates ChatGPT’s broader role in pull request decision-making, including its impact on pull request closures. This includes cases where ChatGPT provides direct code suggestions, as well as instances where it offers conceptual guidance, debugging strategies, or alternative solutions. Building on work by Champa et al. [18] and Chouchen et al. [35] exploring developer interactions with ChatGPT, we examine its potential to streamline decision-making, enhance collaboration, and improve code review processes. By classifying pull request outcomes, our analysis provides empirical insight into ChatGPT’s strengths and limitations in collaborative software development, moving beyond controlled experiments to real-world workflows.
To support this investigation, we extend the DevGPT dataset[37], incorporating additional repositories across diverse programming languages and project contexts. This research expands on an initial study presented at ASE 2024[38], where we introduced PatchTrack and conducted a preliminary analysis of AI-generated patches. Our study significantly extends this work by (1) improving PatchTrack’s classification algorithm, improving the precision of identifying ChatGPT’s influence on pull request outcomes, (2) expanding the dataset to include more closed pull requests, allowing a deeper analysis of AI-assisted pull request rejections, and (3) broadening the scope of research beyond patch application to investigate the reasons pull requests are closed despite AI assistance, such as scope misalignment, patch quality and project constraints. These advancements provide a comprehensive assessment of the role of ChatGPT in pull request workflows, offering insight into the larger impact of AI on collaborative software development and decision-making.
This study provides a foundation for future research in AI-assisted software development by empirically analyzing ChatGPT-assisted pull request workflows, identifying key factors influencing patch integration and rejection, refining automated classification methods, and offering practical recommendations for responsible AI adoption in collaborative coding environments. Our key contributions are as follows.
We analyze ChatGPT-assisted pull request workflows, classifying outcomes to assess AI’s influence on patch integration and rejection, helping developers optimize AI-assisted development.
We enhance PatchTrack’s classification accuracy through benchmark testing and manual validation, improving patch tracking reliability.
We expand the dataset with additional repositories and closed pull requests, enabling a broader analysis of ChatGPT’s role in pull request decisions and guiding AI adoption strategies.
We perform a qualitative analysis of pull request decision-making, identify key factors that affect the integration of AI patches, and provide actionable insights for developers.
We offer practical recommendations for developers, researchers, and educators on responsible AI integration in software engineering workflows.
We release a replication package [39] with datasets, classification tools, and scripts to support reproducibility and future research.
2 Terminology, Problem, and Concrete Examples
In this section, we define key terminology used in this study, outline the classification of ChatGPT-assisted developer interactions, and present a concrete example to illustrate these concepts.
2.1 Terminology
Code snippet: A small, standalone block of code intended for reuse, demonstrating a specific task or programming concept. In this study, “code snippet” refers exclusively to discrete pieces of code that ChatGPT suggests, not including code examples embedded within descriptive text or theoretical guidance.
Diff: The difference between two versions of files, highlighting the changes between an original and a modified file.
Patch: A file containing differences between two versions of one or more files. A patch is a specific type of diff that outlines the changes from one state to another.
Hunk: A contiguous block of changes within a diff or patch, indicating differing lines between two versions of a file [40]. Hunks are identified by headers in the format @@ -s,n +s,n @@, which mark the start and the number of lines affected in the patch.
Self-admitted ChatGPT usage (SACU): Instances where developers explicitly acknowledge, within a pull request discussion or issue thread, that they used ChatGPT to generate, modify, or assist in writing code. This admission is typically evident in pull request comments, commit messages, discussion threads, or issue reports where developers refer to AI-generated suggestions.
2.2 Classifying Developer Interaction with ChatGPT
Figure 1 presents a typical GitHub pull request workflow augmented with ChatGPT interactions, which can originate from the pull request author or a reviewer. Consider a scenario in which a developer, intending to fix a bug, add a feature, or improve an existing one, initiates a change to a file named Foo. Following a pull-based development workflow [41], the developer creates a new branch, commits changes, and eventually merges the pull request into the main repository of the project.
During development, the developer might encounter challenges or seek optimal solutions, prompting them to consult ChatGPT. The responses provided by ChatGPT may be textual explanations or code snippets. In this study, we focus on self-admitted ChatGPT usage (SACU), where developers explicitly acknowledge using ChatGPT within pull request discussions. Based on the self-admitted ChatGPT usage interactions, we classify ChatGPT-assisted pull requests into three possible scenarios:
1.
When the response includes a code snippet, the developer may apply this directly to Foo as a patch, either as provided or after modifications. This scenario is labeled as “Patch Applied (PA)”.
2.
If the code snippets are not applied, or modified beyond recognition (resulting in a semantic clone), Foo may remain unchanged or include changes influenced by the discussed concepts. This scenario is defined as “Patch Not Applied (PN)”.
3.
If the interaction with ChatGPT involves only textual advice without code snippets, similar to the previous scenario, Foo may remain unchanged or may be modified based on the guidance received. This scenario is referred to as “No Existing Patch (NE)”.
Our classification relies on the self-admitted ChatGPT usage, meaning that only cases where developers explicitly state that ChatGPT was used are included in the dataset. This ensures transparency in identifying AI-assisted development while acknowledging that instances of non-disclosed ChatGPT usage remain outside the scope of this study.
2.3 Concrete Example
To put the terminologies and scenarios in perspective, we present a real-world example of “PA” scenario. Concrete examples of “PN” and “NE” cases are discussed later in Section 5. In this example, the developer consulted ChatGPT for assistance with an issue concerning continuous integration. ChatGPT offered a code snippet to address the problem. Figure 2 illustrates a segment of this interaction. The complete dialogue is available here [42] which is included as a reference in pull request #2405 [43] within the faker-js/faker repository on Sept 18, 2023. The analysis of the diff file for pull request (Commit: 62eb120, Hunk: @@ -0,0 +1,31 @@) reveals that the developer integrated the exact code snippet recommended by ChatGPT before merging the pull request on Sept 20, 2023. The lines marked with the symbol in the diff file clearly show the added code. This exemplifies an interesting case, which we categorize as PA. We aim to determine how many similar instances exist in merged pull requests, explore what motivates developers to seek assistance from ChatGPT, and analyze how they use the information obtained from ChatGPT to influence the pull request review and decision-making process.
Beyond the scenarios where the pull request is merged, we also analyze cases where a pull request is closed without merging, despite developers sharing ChatGPT conversation links within the pull request discussions. In these instances, the developer creates a new branch, modifies files, and submits a pull request following the standard workflow. ChatGPT may have assisted with code suggestions, debugging, or conceptual guidance, yet the pull request is ultimately closed due to factors such as scope misalignment, quality concerns, or project constraints. We refer to this scenario as the Closed Case (CL). This study investigates the reasons behind CL, examining how ChatGPT’s involvement influenced closure decisions and whether AI-assisted contributions shaped the final outcome.
3 Study Design
This section presents the research method from data collection, tool design, and quantitative and qualitative analysis. The overarching study design is illustrated in Figure 3.
3.1 Research Questions
RQ1: How are ChatGPT-generated patches (PA, PN, and NE) distributed and utilized in merged pull requests under the self-admitted ChatGPT usage? This RQ investigates the distribution of ChatGPT-generated patches in merged pull requests and explores how developers incorporate them into their workflows. It classifies pull requests into three categories: (i) patches that were applied (PA), (ii) patches that were suggested but not applied (PN), and (iii) cases where ChatGPT did not generate patches (NE). This classification provides insights into the extent of AI-generated patch adoption and its influence on development practices.
RQ2: Why are ChatGPT-suggested patches in conversations not directly integrated (PN) into the pull request, and how do developers use ChatGPT’s suggestions in their workflow under the self-admitted ChatGPT usage? Understanding why developers choose not to integrate ChatGPT-generated patches can provide insights into the challenges and constraints that influence AI-assisted development. Furthermore, examining how developers engage with ChatGPT suggestions can shed light on the broader role of AI in software engineering workflows.
RQ3: Under what circumstances does ChatGPT not generate patches in developer interactions (NE), and how do developers use its responses under the self-admitted ChatGPT usage? This RQ examines cases where ChatGPT does not suggest direct code patches, exploring the reasons behind such interactions. In addition, it investigates how developers incorporate the responses of ChatGPT - whether for conceptual understanding or other non-patch-related tasks - into their development workflows.
RQ4: How does ChatGPT influence closed pull requests, and what factors contribute to these outcomes under the self-admitted ChatGPT usage? This RQ examines the relationship between ChatGPT-assisted contributions and closed pull requests. It aims to explore the factors that lead to pull request closures in cases where ChatGPT-generated content is referenced, providing insights into the broader role of AI-assisted development in collaborative software engineering.
3.2 Dataset Construction and Processing
Figure 3 provides a visual representation of the dataset extraction and processing workflow. The following steps outline the methodology used to identify, extract, and classify ChatGPT interactions within pull requests, ensuring a systematic and reproducible approach to data collection.
3.2.1 Step 1: Identification of merged pull requests containing ChatGPT conversation shared links:
We build on the original DevGPT dataset [37]. This dataset contains nine snapshots collected from GitHub and Hacker News between July 27, 2023, and October 12, 2023, featuring commits, issues, discussions, pull requests, and threads where developers shared links to ChatGPT conversations. Our study focuses only on the merged and closed (unmerged) pull requests included in these snapshots. The initial DevGPT dataset comprises 98 merged pull requests and 39 unmerged (closed) pull requests.
Expanding the DevGPT Dataset. To ensure a comprehensive and up-to-date analysis, we aimed to enlarge the dataset by including more recent instances of merged pull requests and those that are closed without merging but containing ChatGPT conversation links. Using GitHub API v3 [44], we replicated the original data collection methodology with modifications to accommodate data from October 13, 2023, to August 24, 2024. Our search parameters targeted pull requests that specifically included the keyword “https://chat.openai.com/share”. We used a regular expression (“https://chat.openai.com/share/[a-zA-Z0-9-]36”) to precisely extract ChatGPT links from pull request descriptions, commits, or review discussions.
This expanded search resulted in the identification of 187 new merged pull requests and 14 closed and unmerged pull requests that met our criteria. After filtering out inactive or ‘toy’ projects and verifying link functionality, our updated dataset now includes a total of 338 pull requests from 255 unique repositories, providing a robust foundation for analyzing the integration of ChatGPT conversations in software development workflows. As shown in Figure 4, we analyzed repository metrics such as stars, forks, commits, watchers, and language diversity to assess project maturity. The repositories varied significantly in size and engagement. The stars ranged from 0 to 81k (median: 58), and the forks ranged from 0 to 36,655 (median: 31), indicating both niche and widely recognized projects. The commits also showed a wide variation, with some repositories exceeding 36k. Watchers, a measure of visibility, were highly skewed, peaking at over 1.2 million (median: 946), reflecting differences in community interest and activity. These repository attributes have been used in previous studies to measure project diversity and participation [45, 46, 47].
The dataset includes repositories written in 123 distinct programming languages, demonstrating a high level of linguistic diversity. The most frequently used languages are Shell (85.5%), JavaScript (82.0%), HTML (62.0%), TypeScript (50.0%), Python (54.1%), and CSS (54.0%), indicating a strong presence of web development, scripting, and data processing projects. Other widely used languages include Dockerfile (49.0%), Makefile (33.0%), SCSS (22.4%), Java (19.0%), C (18.4%), and C++ (14.5%), reflecting a mix of infrastructure, compiled languages, and styling frameworks.
Although this extensive language diversity suggests that the dataset covers a broad spectrum of software development domains, its distribution is not fully representative of GitHub’s overall ecosystem. Compared to GitHub’s reported trends [48], our dataset overrepresents Shell scripting, web development technologies (HTML, CSS, JavaScript, TypeScript), and infrastructure-related files (Dockerfiles, Makefiles), while underrepresenting widely used languages such as Python and Java. According to GitHub’s Octoverse 2024 report [48], Python has become the most widely used programming language, surpassing JavaScript, and Java, C and C++ also rank highly in the use of the repository. This discrepancy suggests that developers may be more inclined to use ChatGPT for routine scripting, configuration management, and front-end development, where AI-generated suggestions align with established patterns and automation-friendly tasks. In contrast, in domains requiring complex reasoning, such as data science or large-scale backend engineering, developers can rely on specialized frameworks and tools beyond ChatGPT’s current capabilities.
Consequently, while our findings provide strong insights into ChatGPT’s role in pull request workflows within scripting and web development contexts, further research is needed to determine whether similar patterns hold in other software engineering domains. Future work could mitigate this bias by sampling repositories based on GitHub language distribution and investigating AI’s role in diverse development environments, including scientific computing, embedded systems and enterprise software.
3.2.2 Step 2: Extracting pull requests and ChatGPT patches
The pull request and ChatGPT patches are central for analysis in RQ1. Our dataset consists of links to pull requests, ChatGPT links, and patches (if any) extracted from the conversation. The process of extracting patches from ChatGPT conversations was not completely automated, especially in the extended dataset. We noted that OpenAI’s updated terms of service (as of November 14, 2023) no longer permit users to automatically extract data from conversation outputs [49].
To overcome this challenge, we manually converted all the shared links obtained in Step 1 into HTML format and performed scraping on the local files without making HTTP requests to the actual links. pull request patches are extracted using the GitHub API /pulls/ endpoint and saved as “diff” files. At the end of this step, we have two sets of patches for each pull request, namely ChatGPT and pull request patches. This data serves as input for the patch classification step.
3.2.3 Step 3: Patch classification
In this step, and the subsequent two steps, we utilize the PatchTrack patch detection tool, developed by adapting PaReco [50]. While PaReco is designed for the detection of code clones within divergent variant fork software families–identifying missed opportunities, effort duplication, and split cases between mainline and variant branches–it is not directly applicable for detecting ChatGPT patches. In contrast, PatchTrack specifically identifies ChatGPT-shared links, extracts patches, and analyzes prompts. Although PaReco supports exact token matching, its reliance on a Bloom filter data structure leads to higher false positive rates, which can undermine patch identification precision. PatchTrack addresses this limitation by using a temporary array of hash tables for token matching, allowing precise matches with reduced false positives. Hash tables are particularly suitable for exact data retrieval, offering an average-case time complexity of for lookups [51, 52], which contrasts favorably with the lookup time associated with Bloom filters. This design significantly improves PatchTrack’s efficiency and accuracy, making it more effective for exact token matching in diverse software development contexts.
To classify a patch suggested by ChatGPT for a specific pull request, PatchTrack compares it with all other patches in the pull request. PatchTrack performs normalization by removing whitespace, comments, converting characters to lowercase, removing non-ASCII characters, and applying other optimizations tailored to the file types identified by the file extensions of each file. Importantly, these normalization steps are applied to both the ChatGPT and pull request patches being classified. Currently, PatchTrack supports 38 different file types, primarily based on the dataset utilized in this study. These encompass the most commonly used file types within our collected data, enabling the tool to recognize and adapt to the unique syntax and comment styles of each programming language. The architecture of PatchTrack is designed with scalability in mind, allowing easy extension as needs arise or the dataset evolves, ensuring accurate and adaptable patch classification.
Building on methodologies established by tools such as PaReco [50] and ReDeBug [53], PatchTrack employ n-gram tokenization to facilitate precise patch detection. While PaReco and ReDeBug typically utilizes an n-gram value of , our tool uses to accommodate the often smaller size of ChatGPT patches. A detailed comparison of various -gram values is presented in Section 4.1, where we discuss our findings that validate the choice of as optimal for this study.
A patch is classified as PA, PN, NE, “cannot classify” (CC), or “error encountered” (EE). This classification occurs at the Hunk level of each pull request patch. For each hunk, PatchTrack compares the n-gram tokens of the added lines (lines prefixed with ‘’) with those of the ChatGPT codesnippet. If at least one match is found, the patch is classified as PA; otherwise, it is classified as PN. In the absence of a ChatGPT code snippet for a given pull request case, all patches are classified as NE. Currently, PatchTrack supports all file types present in our dataset. Should a file type not yet be supported, it would be classified as CC, indicating areas for future enhancement rather than a limitation. Furthermore, any processing errors encountered during patch analysis lead to an EE classification, typically arising from issues within the tool’s processing functions.
3.2.4 Step 4: Pull request classification
The classification of pull requests in Step 3 follows a straightforward approach. For each class, we aggregate its occurrences within a pull request. If the count of PA is greater than zero, the entire pull request is classified as PA. If there are no instances of PA but there are instances of PN, the pull request is classified as PN. If no ChatGPT code snippets were identified in step 3, both PA and PN counts for this pull request will be zero, leading to an overall classification of NE. The analysis provided by PatchTrack shows how the 645 ChatGPT-generated code snippets were compared with 3,486 pull request patches, resulting in precise classifications and valuable insights into the interaction between developers and AI-generated patches.
3.2.5 Step 5: Compute percentage of ChatGPT patch applied
Since we are performing an exact match in clone detection, we employed the Jaccard containment metric, defined as:
Where and represent the token streams from ChatGPT and GitHub patches, respectively. This metric calculates the proportion of common code by dividing the number of tokens at the intersection of and by the total number of tokens in the ChatGPT suggested patch. This metric was adopted from the original Redebug paper [53]. The result shows the percentage of ChatGPT’s code integrated into the pull request by the developer.
3.2.6 Step 6: Qualitative analysis
To address RQ2, RQ3, and RQ4, we employed a rigorous qualitative analysis methodology. This involved summarizing ChatGPT developer conversations and GitHub pull requests using the Framework Method [54], followed by thematic extraction using the card sorting technique [55]. This process was carried out collaboratively by 2 Ph.D. students and 1 Professor from the same lab.
Step 6.1: Summarization using the Framework Method:
We collected relevant ChatGPT–developer conversations and GitHub pull requests categorized as PN and NE, organized into a spreadsheet. In the pull requests, we specifically focused on the sections where the ChatGPT link was shared, examining the code and developer interactions a few steps before and after the link to ensure relevance to the ChatGPT conversation, as the pull requests might contain unrelated content.
Each team member familiarized themselves with the data, and we developed an initial coding framework based on our research questions, focusing on elements such as the nature of the advice given by ChatGPT, the type of challenge encountered by the developer, and the reasons for non-integration of patches or absence of patch suggestions.
Through iterative discussions, we refined this framework and systematically applied it to the data, creating summaries for each interaction. These summaries captured key information such as the developer’s problem, the advice provided by ChatGPT, how the developer applied or did not apply this advice, and the outcomes of the interaction.
To enhance the quality of our summaries, we leveraged ChatGPT version 4.0, providing it with organized ChatGPT developer conversations and the corresponding pull requests for each instance. We compared ChatGPT’s summaries with our manual ones, finding ChatGPT’s results often superior in English context and clarity. However, in instances where ChatGPT produced hallucinations, we resorted to our manually constructed summaries. For inter-coder reliability, each team member independently summarized a subset of cases, and we reached a consensus on the final summaries.
Step 6.2: Card Sorting for Theme Extraction:
Using the summaries created in the Framework Method, we prepared detailed cards for each interaction. Each team member individually sorted the summary cards into categories based on perceived similarities and emerging themes, grouping the cards into thematic clusters. After individual sorting, we conducted a group discussion to compare and consolidate the categories, refining and agreeing on the final set of themes. We then performed a thematic analysis on the agreed categories to identify and define the overarching themes, reviewing them for coherence, refining them to ensure that they accurately represented the data, and naming them appropriately.
4 Results & Discussion - RQ1
How are ChatGPT-generated patches (PA, PN, and NE) distributed and utilized in merged pull requests under self-admitted ChatGPT usage?
This research question quantifies the distribution of pull requests classified as PA, PN, and NE within our 285-case dataset.
To support this classification, we evaluated the tokenization strategy used to compare ChatGPT-suggested code with pull request patches. Since ChatGPT-generated snippets are often brief, selecting an appropriate n-gram size is critical to capturing exact matches while minimizing false positives. We conducted a comparative analysis across n-gram values , and , validating results with both automated metrics and manual inspection.
4.1 Token Matching Methodology
Our evaluation showed that consistently achieved the best performance, accuracy (97.5%), precision (96.5%), recall (96.7%), and F1-score (96.6%), particularly for detecting short and exact matches. Larger values reduced recall and did not capture concise code fragments. Manual validation (Cohen’s Kappa = 0.85) further confirmed that provided the most reliable alignment between the suggestions generated by ChatGPT and the patches applied by the developer.
We also examined how token size influenced patch classification outcomes. Smaller n-grams (n = 1) identified more patch applied (PA) cases and fewer patch not applied (PN) cases, whereas larger n-grams were more sensitive to formatting and syntactic variations, leading to under-detection. The complete results are provided in the appendix A.
Misclassification Analysis
We identified and manually analyzed 7 misclassified cases (4 patch not applied PN and 3 patch applied PA), to better understand PatchTrack’s limitations under the setting. Two key factors contributed:
1.
Inaccessible Patches via GitHub API: GitHub does not expose file diffs exceeding 400 lines via its API. As a result, large patches were omitted during extraction. In three cases, ChatGPT-generated code was present in these omitted sections, causing false PN classifications (e.g., PR #2232 [56]).
2.
Single-Line Matches: While the tool correctly flags one-line matches, this sometimes led to misclassifying a PN case as PA. These borderline cases highlight a trade-off between sensitivity and semantic depth.
Overall, while PatchTrack effectively identifies directly applied patches, it does not capture structural or semantic transformations. Because it relies on token-level matching rather than higher-order semantic analysis (e.g., Type-4 clone detection), cases with significant refactoring.
4.2 Patch Classification
We analyze 645 ChatGPT-generated code snippets and 3,486 patches in 285 pull requests. To systematically classify these snippets, we employed PatchTrack, which compares each ChatGPT-generated snippet with all patches in its corresponding pull request. This approach enabled us to categorize the snippets into predefined integration groups. Through a series of experiments, we found that using an n-gram value of achieved the highest accuracy, precision, recall, and F1 score compared to other sizes of n-grams. This choice was guided by the concise nature of the ChatGPT-generated patches, where exact token matching improves detection accuracy while minimizing the risk of overlooking minor but significant similarities (see Table 2–appendix A and Section 4.1 for a detailed analysis). Figure 5 summarizes the classification results, illustrating the distribution of pull requests in PA, PN, and NE.
4.2.1 Percentage of ChatGPT Patches Applied
To measure the integration of suggested ChatGPT patches into pull requests, we used Jaccard’s containment ratio 1, which quantifies the overlap between suggested patches and the final code integrated into pull requests. Figure 6 illustrates the distribution of integration percentages for the classification PA. The median integration rate is 25%, indicating that, on average, a quarter of the ChatGPT-suggested code is adopted in merged pull requests.
Developers integrate ChatGPT patches with varying degrees of modification, assessing their relevance based on project needs. While some adopt AI-generated patches wholesale, others incorporate only selective portions. This variation reflects how developers balance AI-assisted suggestions with project-specific constraints such as maintainability, readability, and adherence to coding standards. The distribution of integration percentages, as shown in Figure 6, highlights this diversity in adoption, with some pull requests fully incorporating ChatGPT-generated code, while others retain only fragments. The presence of outliers in the distribution underscores the exception rather than the norm: while a few developers fully adopt ChatGPT-generated patches, the majority incorporate only parts of the suggestions, often modifying them extensively to fit project-specific requirements.
The median integration rate of 25% (Q2) suggests that full adoption of ChatGPT-generated patches is uncommon, with most pull requests retaining only a fraction of the AI-suggested code. The third quartile (Q3) at 38% further reinforces that even in cases with higher integration, significant modifications are often required. These findings indicate that developers frequently treat the ChatGPT output as a starting point rather than a final implementation, selectively incorporating code that aligns with project needs.
To better understand this variation, we selected four pull requests representing different levels of integration, with one pull request chosen from each quartile: (i) low (0–25%], where only minor elements of the suggested patch were retained, often requiring substantial modifications or alternative implementations; (ii) lower-median (25–50%], where developers incorporated portions of the patch but refined it significantly; (iii) upper-median (50–75%], where a larger portion of the patch was adopted with some modifications; and (iv) high (75–100%], where the AI-generated patch was largely accepted with minimal changes. Each quartile follows the standard convention where the lower bound is excluded, and the upper bound is included. These cases were selected through stratified random sampling to ensure balanced representation across integration levels. By analyzing these strata, we illustrate real-world scenarios of ChatGPT-assisted patch integration, shedding light on why some patches are fully applied, while others are selectively merged or heavily revised.
Example 1: Low Patch Integration (0–25%] In pull request #185[57] of the pokt-network/poktroll repository, the developer asked ChatGPT111Example 1 link to ChatGPT conversation – Accessed: 2025-02-17 if Go had a built-in function to remove leading spaces from multiline strings while maintaining code readability. ChatGPT suggested an approach using strings.Split, strings.TrimLeft, and strings.Join to process each line and remove leading spaces. The developer then requested a refined solution that encapsulated this logic into a reusable helper function, trimString, to improve readability. ChatGPT provided a full Go program, but only its helper function was integrated. Of the suggested 31 LOC, just 7 LOC (22.6%) were adopted. The minimal adoption reflects the developer’s selective integration of only the core functionality needed, as the existing project structure did not require a complete rewrite.
Example 2: Lower-median Patch Integration (25% - 50%] pull request #7123 [58] in the Mudlet/Mudlet repository addressed a crash that occurred when double-clicking on a word in TTextEdit. The issue stemmed from an out-of-bounds access in a function responsible for expanding text selection, leading to an assertion failure in a QString. After encountering a stack trace pointing to the faulty function, the developer consulted ChatGPT222Example 2 link to ChatGPT conversation – Accessed: 2025-02-17, which identified the issue and suggested a structured fix: (i) Implement boundary checks for yind, (ii) Break loops when indices become invalid, and (iii) Refactor selection logic to improve robustness. The initial ChatGPT-generated patch (39 LOC) effectively prevented the crash by halting execution before an invalid index was accessed. However, only 18 LOC (46.2%) were ultimately integrated after a reviewer suggested negating conditions to reduce indentation and improve readability. The extent of the modifications suggests that while the ChatGPT recommendations provided a strong starting point, the developers refined the patch to better align with the project constraints and coding practices. This case illustrates how AI-assisted debugging can be useful but requires human oversight and adjustments to ensure code quality and maintainability.
Example 3: Upper-median Patch Integration (50% - 75%] pull request #279 [59] in the nylas/nylas-python repository introduced support for a new “free-busy” API endpoint in the Nylas SDK, enabling users to check calendar availability for specific email addresses. A key challenge in this implementation was designing a suitable data model to represent the API response, which could include both availability data and error messages. To address this, the developer consulted ChatGPT333Example 3 link to ChatGPT conversation – Accessed: 2025-02-17 for guidance on structuring Python dataclasses for the JSON response. ChatGPT suggested: (i) defining separate data models for distinct response components, (ii) utilizing Python’s dataclass module for maintainability, and (iii) employing Union types to handle multiple object types within a single list. The proposed patch consisted of 25 lines of code (LOC), of which 15 LOC (60%) were integrated. The primary modification involved renaming the error class to freeBusyError for improved clarity, based on reviewer feedback. This case highlights how AI-generated code can closely align with project requirements, yet still requires refinements guided by domain expertise to enhance readability and maintainability.
Example 4: High Patch Integration (75–100%] In pull request #2059 [60] of the alshedivat/al-folio repository, a developer extended a Jekyll plugin to support fetching posts from a manually specified list of URLs in addition to an RSS feed. Initially, the plugin retrieved posts only via RSS, but the developer aimed to enhance its functionality by incorporating direct URL input. To achieve this, they consulted ChatGPT444Example 4 link to ChatGPT conversation – Accessed: 2025-02-17 to modify the Ruby-based Jekyll plugin, focusing on three key aspects: (i) metadata extraction, (ii) correct handling of published dates, and (iii) improved logging readability. ChatGPT contributed by refactoring the code to introduce modular functions, ensuring robust date handling with Time objects, improving log clarity through puts, and enhancing maintainability by extracting date parsing into a dedicated parse_published_date function. The patch consisted of 19 lines of code (LOC), all of which were integrated as-is (100%), as it aligned perfectly with the project’s needs, requiring no modifications. The developer’s enthusiasm, expressed at the end of their interaction with ChatGPT (“Awesome! I have a feeling we’ll write a lot of code together from now on ;)”), underscores how ChatGPT’s response was fully applicable without requiring further adjustments.
4.3 Discussion and Implications - RQ1
Our findings indicate that while ChatGPT-generated patches provide value, full adoption remains rare, with a median integration rate of 25%. Developers often treat AI-generated suggestions as a starting point rather than a final implementation, selectively integrating elements that align with project-specific constraints. The third quartile (38%) further reinforces that even in cases with higher integration, significant modifications are frequently required. Integration patterns reveal different adoption behaviors: In low integration cases (0-25%), developers retained only minor portions of AI-generated code due to inefficiencies, redundancy, or misalignment with project conventions. Moderate integration cases (25–75%) show that developers adopted substantial sections but refined them to fit architectural guidelines, while high integration cases (75–100%) occurred when ChatGPT-generated patches closely matched project needs, requiring little or no modification.
The predominance of low- and moderate-integration cases highlights key challenges in AI-assisted development, such as misalignment with project coding styles, maintainability concerns, and trust issues. This underscores the need for AI tools to incorporate better contextual awareness, generating suggestions that require minimal refinement. Although ChatGPT reduces the initial effort to write code, its effectiveness depends on the developer’s ability to evaluate, adapt, and refine AI-generated patches. At the same time, high-integration cases demonstrate that well-aligned AI-generated patches can streamline the pull request process, reinforcing AI’s potential in accelerating development workflows.
In particular, our study focuses on self-admitted ChatGPT usage (SACU), which means that all the cases analyzed involve developers explicitly stating their use of ChatGPT within pull request discussions. This ensures transparency in identifying AI-assisted contributions, but also introduces a limitation: we do not capture instances where ChatGPT was used privately or without disclosure. The question of whether undisclosed ChatGPT usage follows a different pattern of patch integration, which could influence the observed distribution of patch applied (PA), patch not applied (PN), and no existing patch (NE) cases remains open.
These observations align with the longstanding vision of the “Programmer’s Apprentice” [15], which imagined AI as a reasoning collaborator in software development workflows. In this vision, AI acts not as an autonomous agent, but as a junior partner that supports design exploration, debugging, and solution refinement while leaving final decisions to the developer. Our findings reflect this paradigm: developers often treat the ChatGPT output as a starting point, adapt it based on the project-specific context, and integrate it selectively. This underscores the role of ChatGPT as a decision-support tool rather than a turn-key solution in pull request workflows.
Although our analysis quantified the extent to which ChatGPT-generated patches were integrated, it did not perform a qualitative examination of why some patches were applied more successfully than others. Future work could investigate the characteristics of applied patches, such as query formulation, patch complexity, and developer perception of AI-generated solutions. Understanding these aspects could further refine AI-assisted coding and improve AI effectiveness in collaborative development. Furthermore, this discussion sets the groundwork for RQ2 (why patches are not applied), RQ3 (why no patches are suggested), and RQ4 (why pull requests are closed despite the participation of ChatGPT), facilitating a deeper examination of the role of AI in pull request decision-making.
5 Results and Discussion – RQ2, RQ3 & RQ4
Recall that to address RQ2, RQ3, and RQ4, in Section 3.2–Step 6, we employed a rigorous qualitative analysis methodology involving summarizing ChatGPT-developer conversations and GitHub pull requests, followed by thematic extraction using a card sorting technique.
For brevity, we present one representative instance per theme in the results sections. Additional examples and detailed summaries are available in the replication package [39]. The full set of data, including all summaries, identified themes, pull request links, and corresponding ChatGPT-developer conversation links, is provided in the replication package to support reproducibility and further analysis. Each instance is labeled with a unique identifier: PN-X (X being the instance number) for patches not applied, NE-X for cases where no patches were suggested and CL-X for closed pull requests, to facilitate cross-referencing within the replication package [39].
[…]
5.1 Results and Discussion - RQ2
Why are ChatGPT-suggested patches in conversations not directly integrated (PN) into the pull request, and how do developers use ChatGPT’s suggestions in their workflow under the self-admitted ChatGPT usage?
In our analysis, the tool initially classified 63 patch not applied (PN) instances. However, as detailed in Section 4.1, a manual review was performed on a sample of 55 PN cases to validate the precision of the classification. This review identified 3 false positives that were incorrectly classified as PN and 4 false negatives from patch applied (PA) that should have been included in PN. After these corrections, we obtained 56 valid PN instances for thematic analysis. The analysis of these 56 PN instances revealed six distinct themes, as illustrated in Figure 7.
Adaptation and tailored solutions: This emerged as the most prevalent theme, with 17 of 56 cases showing that developers adapted the conceptual insights of ChatGPT to fit specific project requirements. For example, in pull request #12 [61] – PN-19, a developer sought a regular expression from ChatGPT to match the ULID format in the laravel-json-api/core project. The developer initially considered ChatGPT’s regex: “Like I said, I didn’t find an ’official’ regex for the ULID. ChatGPT gave me this: […]”. However, rather than directly applying the AI-generated solution, the developer aligned the implementation with Laravel’s native method for handling ULIDs. This decision was influenced by the comment of a reviewer: “I think we should match what Laravel does” for route regexes for ULIDs. Looking at this docs: […]. The author acknowledged the feedback: “[…] good idea […]. Change made!” and updated the code to ensure consistency within the Laravel ecosystem. This example highlights how developers use ChatGPT as a starting point, but ultimately refine AI-generated suggestions to conform to established coding conventions and project-specific constraints.
Technical limitation and preferences: This was the second most prominent theme, with 9 of 56 instances demonstrating how ChatGPT’s patches were not applied due to technical constraints or developer preferences for alternative solutions. In these cases, developers faced trade-offs in implementation strategies, either due to limitations in the existing system or subjective choices regarding the best approach to a problem. A representative example is pull request #171 [62] within the ory/elements project (PN-11), where developers attempted to disable a button when submitting the form. A reviewer sought ChatGPT’s input on a possible fix, asking: “Wouldn’t the fix just be to change click to submit? […]” However, another reviewer pointed out a critical issue with this approach: “This doesn’t actually work. The reason is that the click event fires before the submit event […] but IMO we should just get this right […]” The pull request author ultimately acknowledged that resolving the issue properly would require additional time and effort, stating: “[…] just don’t want to spend time on getting this right, right now […]” Upon further examination, we found that the proposed solution was reverted because disabling the button caused unintended side effects, such as stopping event propagation. Given the minor impact of the issue, the team decided to focus on more pressing updates rather than implementing a workaround.
Methodological Guidance: We identified 8 of the 56 instances in which ChatGPT influenced developers’ decision-making strategies by suggesting optimizations or best practices rather than providing direct fixes. For example, in pull request #5063 [63] within the darklang/dark project, a developer attempted to resolve an issue related to function composition. A reviewer, finding the initial implementation overly complex, consulted ChatGPT, which suggested a more concise approach. The reviewer commented: “While this works, it feels like overkill. I asked ChatGPT about how best to do this, and it suggested using (int)c. That seems to me like it would work. ChatGPT-link[…]”. This led the pull request author to reassess the solution, ultimately adopting a simplified version influenced by ChatGPT’s recommendation.
Guidance on Functionality-Related Decisions: Developers often seek to modify or enhance specific functionality in software projects, often consulting ChatGPT for insights on best practices or alternative implementations. This theme captures instances where ChatGPT’s recommendations influenced functionality-related decisions, even when no direct patches were applied. From our dataset, we identified 8 of the 56 instances in this category. A representative example is pull request #522 [64] within the sveltejs/learn.svelte.dev repository (PN-13). In this case, the pull request author sought guidance on the correct ARIA role for an interactive SVG acting as a pointer in a Svelte application. The conversation focused on whether to use “role=none” or “role=presentation.” ChatGPT provided contextual insights on ARIA accessibility, suggesting that “role=presentation” would be more appropriate based on the functionality of the SVGs. The pull request description explicitly references ChatGPT’s recommendation: “presentation” might be better suited here according to ChatGPT: […]. The integrator acknowledged the contribution with a brief confirmation: “Thank you!” and merged the pull request. This case illustrates how ChatGPT’s suggestions can shape design decisions and accessibility improvements, with maintainers providing the final validation before integration.
Validation and Refinement of Implementation: Developers often use ChatGPT to validate their reasoning and refine their decision-making process rather than applying AI-generated patches directly. In 8 of 56 instances, the interactions focused on seeking conceptual clarification or evaluating alternative implementations before finalizing a solution. This indicates that AI tools play a role not only in code generation, but also in developer cognition and reasoning during pull request workflows. A representative example is pull request #5058 [65] – PN-28, where a developer was considering implementing the List.filter function using recursion versus List.fold in a functional programming context. A reviewer commented: “I was thinking about whether recursive implementations were more appropriate here, since this can also be implemented with List.fold. […] You might enjoy what ChatGPT told me about it: […]” This prompted further discussion about theoretical best practices in functional programming, highlighting how ChatGPT is used as an advisory tool rather than a direct source of patches. Ultimately, rather than replacing developer decision-making, ChatGPT functioned as a catalyst for discussion and refinement within the pull request workflow.
Project Organization and Documentation Refinement: Beyond code modifications, developers frequently use ChatGPT to help improve project organization, improve documentation, and optimize workflow structures. This theme captures instances where ChatGPT’s suggestions influenced non-code contributions, such as file management and documentation strategies, rather than direct patch implementations. From our dataset, we identified 6 of the 56 instances in which ChatGPT was consulted to streamline internal project structures. A representative case is pull request #37 [66] within the Bananapus/nana-core project (PN-2), where the developer sought an efficient way to manage file renaming within a pull request. The ChatGPT prompt specified the issue: “in a GitHub pull request, I have renamed a file. In the diff viewer, it shows one file deleted and another added, but I instead want to show the file as renamed.” ChatGPT suggested a structured file management approach to ensure that Git detected the renaming action instead of treating it as deletion and addition. The developer applied this suggestion, improving repository maintainability and commit traceability. No code patch was applied since the pull request focused on internal organization rather than functional modifications.
[…]
5.1.1 Implications for RQ2
Our findings indicate that developers rarely apply ChatGPT-suggested patches verbatim. Instead, they critically assess AI-generated suggestions, modifying them to align with project requirements, maintainability constraints, and workflow considerations. Even when patches are not directly integrated, ChatGPT contributes to software development by shaping decision-making, guiding optimizations, and supporting documentation improvements. These insights have direct implications for both practitioners and researchers in understanding how AI-assisted workflows function when patches are not applied. Recent empirical evaluations have highlighted fundamental limitations in LLMs’ ability to understand, reason about, and modify source code [67, 68, 69]. These limitations, including difficulties in semantic comprehension and contextual adaptation, may partly explain why ChatGPT-suggested patches often required developer refinement or were ultimately not applied in their original form. Our findings reinforce the need for critical evaluation of AI-generated suggestions within collaborative development workflows.
[…]
Developers use ChatGPT’s suggestions as conceptual starting points rather than direct solutions, often modifying or rejecting them based on project-specific requirements. This underscores the need for structured validation when using AI-assisted patches. Since ChatGPT-generated code may not always be directly compatible with existing architectures or coding styles, developers should critically assess AI suggestions within the context of system dependencies, maintainability goals, and pull request review feedback before applying them. This finding suggests that AI-assisted patch adoption can be improved by establishing clearer internal guidelines on when and how to incorporate ChatGPT’s code into ongoing development workflows.
Beyond direct code generation, ChatGPT influences how developers refine and document their implementation strategies. Our results show that even when ChatGPT’s patches are not applied, developers use AI-generated insights to improve documentation clarity, enforce project-specific conventions, and inform architectural decisions. This suggests that teams can benefit from integrating AI tools into pull request discussions, even when patches are not used directly, by treating them as a mechanism to refine implementation details and improve project documentation.
[…]
These findings highlight a crucial research gap to understand how AI-generated suggestions are evaluated, modified, or dismissed prior to pull request integration. Future research should analyze how developers iteratively refine AI-assisted patches, distinguishing between modifications that preserve AI-generated logic versus those that entirely rework AI suggestions. Studying decision patterns, such as why some ChatGPT-suggested patches are partially modified while others are rejected, could provide insights into when AI-generated solutions are most effective.
In addition, improving AI-generated patches requires more context-aware recommendation models. Current AI systems lack fine-grained awareness of project-specific constraints, repository structures, and architectural dependencies, leading to suggestions that are syntactically valid but not semantically useful. Future research could explore adaptive AI models that learn from previous pull request discussions and historical development patterns, ensuring that AI-suggested patches align with specific project norms and are more seamlessly integrated into pull request workflows. This refinement would help bridge the gap between AI-assisted coding and real-world software engineering practices.
5.2 Results & Discussion - RQ3
Under what circumstances does ChatGPT not generate patches in developer interactions (no existing patch-NE), and how do developers use its responses under the self-admitted ChatGPT usage?
From the generated summaries of the 84 NE instances, the employed card sorting analysis yielded four distinct themes, as illustrated in Figure 8. The identified themes are: Conceptual Guidance & Theoretical Advice, Documentation, Communication & Translation, Education and Knowledge Sharing, and Debugging & Optimization Strategies.
Conceptual Guidance & Theoretical Advice: This was the most prevalent theme, with 44 of 84 instances where ChatGPT did not suggest patches but instead guided pull request decision-making. Developers refined code based on programming concepts, design principles and best practices, improving readability, maintainability, and overall design. For example, in pull request #1061 [70] within the codecrafters-io/frontend project (NE-36), a reviewer used ChatGPT to critique a variable’s naming convention, stating: “Bad naming, this should be something like selectedPricingFrequency. Reasoning: [ChatGPT link]” The pull request author acknowledged the issue and noted: “[…] we don’t seem to be getting better with naming […] talk to ChatGPT about it […] we need to pay extra attention to it.” Originally, the variable name pricingFrequencyClicked implied an action (clicking) rather than a state (selected frequency). After reviewing ChatGPT’s feedback, the developer renamed it to selectedPricingFrequency, improving clarity. While no code snippet was suggested, ChatGPT’s advice directly influenced the pull request decision, refining naming conventions. The absence of a patch did not hinder decision-making but encouraged refinement. Reviewers used ChatGPT’s insights to justify changes, while maintainers ensured clarity before merging. These interactions shifted pull request discussions from issue resolution to proactive quality improvements, reinforcing AI’s role in better development practices.
Documentation, Communication & Translation: In 23 of the 84 instances, the discussions centered on refining the documentation, improving communication and ensuring accurate localization. Although no code snippets were generated, ChatGPT’s responses helped developers refine explanations, adjust terminology, and align documentation with project standards before merging changes. A key example is pull request #2230 [71] from the faker-js/faker project (NE-8), where a reviewer suggested re-writing a comment to be more positive and inclusive. The original comment stated: “It would not be a good idea to have an English-only fallback list, because the Esperanto community is very diverse and international.” To improve the phrasing, the reviewer consulted ChatGPT, who suggested multiple re-wordings. The final version was collaboratively adjusted and implemented, with the pull request author noting: “Thank you for your comment, I used this suggestion!” Although ChatGPT did not provide a code patch, its response influenced pull request decision-making by shaping documentation revisions before merging. AI-assisted refinements helped drive consensus in pull request reviews, leading to clearer and more inclusive communication. Although ChatGPT did not generate a code patch, its response influenced pull request decision-making by reinspecting documentation before merging. AI-assisted suggestions helped reviewers justify edits and maintainers ensure clarity, shifting pull request discussions from minor wording adjustments to broader quality improvements.
Education and Knowledge Sharing: This theme captures 9 of the 84 instances in which developers used ChatGPT to clarify programming concepts, system behaviors, or language-specific features. Instead of providing a direct patch, ChatGPT’s explanations helped developers make informed adjustments in pull requests by improving their understanding of technical constraints. A notable example is pull request #2415 [72] from digitalbitbox/bitbox-wallet-app (NE-57), where a developer sought to determine the location of the Homebrew binary installation on a Mac M1 and how different commands influenced the installation paths: “On a Mac M1, what’s the location of the brew binary?” ChatGPT clarified Homebrew installation differences, helping the developer determine the correct binary path. The pull request discussion validated this approach, with the reviewer finding the insight interesting and raising concerns about system-wide effects, such as “replace their normal brew […] all future installs […] wrong architecture.” While no patch was generated, ChatGPT’s response directly influenced the pull request by helping the developer refine installation logic. Understanding how Homebrew installs based on system architecture enabled a more informed decision on script modifications. AI-assisted clarification also guided pull request discussions, leading to further validation and refinement before merging.
Debugging and Optimization Strategies: We identified 8 of the 84 instances where ChatGPT assisted with debugging, performance optimization, and algorithm refinement, providing strategic insights without generating code snippets. An illustrative example is pull request #5068 [73] from the darklang/dark project (NE-2), where a reviewer sought an alternative to improve the performance of using ‘vector‘ in C++, recalling: “I remember you mentioning that the performance of pushback isn’t great. Do you recommend a different approach?” The pull request author consulted ChatGPT for performance evaluation and alternative strategies. After reviewing the AI-generated insights, the author proposed a new approach: “[…] appending it to the front and then reversing should do it […]” The developer switched to using push instead of pushBack, improving performance since inserting at the front typically has time complexity, whereas pushBack could lead to . This discussion shaped the pull request decision, as ChatGPT provided optimization strategies rather than a direct coding solution. Although no patch was generated, ChatGPT influenced pull request decisions by helping the developer assess trade-offs and select a more efficient approach. AI-assisted insights guided the pull request review, enabling an informed optimization decision before merging.
[…]
5.2.1 Implications for RQ3
Our findings highlight that ChatGPT frequently provides non-patch responses that still influence pull request decision-making. Developers do not always require direct patches; instead, they leverage AI-generated feedback for conceptual improvements, documentation refinements, knowledge expansion, and debugging strategies. These insights have important implications for both practitioners and researchers.
[…]
ChatGPT’s non-patch responses help developers refine their code by offering structured feedback on best practices, naming conventions, and maintainability. AI-generated insights can support code reviews, enabling teams to assess design choices and enforce coding standards. In documentation workflows, ChatGPT improves clarity and consistency, allowing teams to incorporate AI-driven language refinements into their review processes. This improves the readability of the documentation and ensures technical accuracy before merging. For debugging and optimization, ChatGPT provides strategic recommendations rather than direct patches, helping developers identify bottlenecks, assess trade-offs, and refine performance strategies. Developers can use these insights to justify pull request decisions and explore alternative implementations before committing changes. By integrating ChatGPT into pull request workflows, developers can enhance collaborative decision-making, using AI not just for code generation but as a real-time knowledge assistant that strengthens software development practices.
[…]
Future research should explore how AI-generated non-patch responses impact long-term software design decisions. Investigating whether repeated exposure to AI-driven conceptual guidance influences coding best practices and architectural choices could provide insights into AI’s role in shaping developer decision-making over time. More studies could examine how AI explanations affect debugging efficiency at different experience levels. Research could assess whether AI-driven insights help junior developers bridge knowledge gaps or provide novel perspectives for experienced engineers in complex problem-solving scenarios. Furthermore, analyzing when and why developers trust AI-generated theoretical advice without direct patches could inform the design of AI systems that better align with developer expectations in collaborative software development.
5.3 Results & Discussion - RQ4
How does ChatGPT influence closed pull requests, and what factors contribute to these outcomes under the self-admitted ChatGPT usage? To address this research question, we conducted an in-depth analysis of 47 closed pull requests that involved interactions between developers and ChatGPT. The qualitative analysis identified six key themes, as illustrated in Figure 9. The most prevalent reason for pull request closure was attributed to Scope and Quality Issues, which accounted for 31.9% of the cases. Other reasons included procedural and technical adjustments (19.1%), experimental solutions (17.0%), administrative or policy issues (10.6%), duplicate solutions (10.6%) and cases where no specific reason (10.6%) was documented.
Scope and Quality Issues This was the leading cause of pull request closures, accounting for 15 of 47 cases. Many pull requests were rejected due to quality concerns, workflow misalignment, or maintainability issues. Although ChatGPT’s optimizations were useful, developers struggled to integrate them into long-term architectural plans. In pull request #428 [74] of the gemini-hlsw/scheduler project (CL-31), ChatGPT suggested replacing inefficient list-based membership checks with a set-based approach and explained Python’s frozenset.issuperset() behavior along with __contains__ and __hash__. Although this improved efficiency, it did not resolve architectural issues related to object consistency during unpickling. Reviewers referred to ChatGPT’s responses to validate assumptions, but ultimately favored a fundamental architectural change over a band-aid fix. They emphasized the need for a long-term solution, stating: “This needs a proper fix rather than a band-aid solution; future developers will face the same issue if left unresolved […]” and “[…] A generator should be used here. If we proceed with this approach (which raises red flags), it must be carefully revised […]”. ChatGPT’s optimization was valid, but the main issue was ensuring object consistency during unpickling. Rather than modifying the membership test, reviewers restructured unpickling by integrating it with ResourceManager and relocating it to the lucupy library. This case shows that, while ChatGPT-assisted optimizations can enhance performance, they may not align with architectural goals. The reviewers found that its insights were useful for validation, but prioritized structural changes for long-term sustainability.
Procedural and Technical Adjustments: This theme encompasses pull requests requiring refinements or procedural modifications before integration. 9 of the 47 cases fell into this category, often involving technical restructuring, refactoring, or adherence to project workflows. A representative case is pull request #534 [75] from theosanderson/taxonium - CL-47, which aimed to introduce a map visualization displaying sample origins in a phylogenetic tree. During the discussions, a reviewer explored the implementations using ChatGPT, which suggested leveraging DeckGL for rendering pie charts. The reviewer noted: “[…]This was ChatGPT’s idea for that in DeckGL[…]”. The pull request author acknowledged the approach’s feasibility, stating: “[…]Neat! I think that approach could work. Let me see if I can get something working in a feature branch[…]”. Although viable, maintainers identified the need for significant back-end modifications, including restructuring metadata handling and creating new API routes. Due to these broader architectural changes, the pull request was closed in favor of continued development within the add-map-view feature branch. This case highlights how pull requests requiring substantial refinements are often redirected rather than merged immediately, aligning with trends in this category. ChatGPT played an exploratory role in shaping the implementation strategy, but the final decisions were guided by project-specific requirements.
Experimental Solutions: 8 of the 47 closed pull requests were later replaced by more refined implementations. These pull requests were often experimental, serving as proof-of-concept solutions that were ultimately superseded by optimized alternatives. In some cases, developers identified existing solutions that made ChatGPT-assisted contributions unnecessary for direct integration. For example, pull requests #2214 [76] and #2212 [77] in the open-learning-exchange/myplanet repository (CL-29 and CL-36) introduced ChatGPT-generated French and Arabic translations for UI strings, addressing issues #2216555https://github.com/open-learning-exchange/myplanet/issues/2216 and #2219666https://github.com/open-learning-exchange/myplanet/issues/2219. However, both pull requests were later closed, with their contributions consolidated into pull request #2229777https://github.com/open-learning-exchange/myplanet/pull/2229, which was merged. This closure pattern aligns with the common practice of refining and merging multiple contributions into a single comprehensive pull request. Although ChatGPT facilitated the initial translation drafts, final refinements were likely developer-led to ensure linguistic accuracy and adherence to project standards. This underscores how ChatGPT-assisted contributions often act as stepping stones toward more polished, project-aligned solutions rather than directly leading to merged changes.
Administrative or Policy Issues: five pull requests were closed due to administrative hurdles, including issues of the signing of the Contributor License Agreement (CLA), policy violations and housekeeping rules such as handling stale pull requests. A notable example is pull request #468 [78] in the labdao/plex project – CL-5, where ChatGPT assisted in the development of a well-structured README.md for Gnina on the PLEX platform. Despite the quality of the documentation, the pull request was eventually closed after being flagged as stale. This finding highlights that ChatGPT’s assistance does not always lead to pull request integration, as non-technical factors can prevent contributions from being merged, aligning with the goal of RQ to understand the role of AI in pull request closures.
Duplicate Solutions: pull requests in this category were closed because they duplicated existing efforts in the codebase or overlapped with other pull requests addressing the same issue. This includes cases where developers unknowingly submitted redundant pull requests or discovered an existing solution after submission. We identified five instances that reflect this theme. For example, in pull request #3792[79] of the plausible/analytics project - CL-20, the developer attempted to optimize how imported views are handled per visit by refining calculations in the stats/base.ex module. The primary concern was whether repeated expressions in queries to ClickHouse led to redundant computations. ChatGPT was consulted to verify if ClickHouse recalculates identical expressions multiple times or reuses results for optimization. Based on the response of ChatGPT, the developer confirmed that ClickHouse avoids redundant calculations, improving the efficiency of the query. However, pull request #3792 was later closed in favor of pull request #3830, which provided a cleaner and more efficient solution. The author of pull request #3792 acknowledged the duplication, commenting: “[…] Closing in favor of #3830.” and later stating: “[….]Uff, sorry for this […]. I had absolutely no intention of redoing your work **emoji** Unexpected coincidence **emoji**”. This case highlights that while ChatGPT helped validate an approach, it did not prevent redundant work, as broader project awareness and coordination among contributors played a more decisive role in the pull request’s closure.
No Reason: This theme captures 5 instances where the reason for pull requests closure was either not documented or remains ambiguous. Without clear explanations or comments from the project maintainers, the rationale for closing these pull requests is left open to interpretation. For instance, pull request #2213[80] proposes the addition of Somali translations for various user interface (UI) strings within the open-learning-exchange/myplanet project - CL-9. The pull request aims to address issue #2218 by incorporating localized Somali translations for menu items, error messages, buttons, and other textual elements within the application. The developer requested ChatGPT to translate UI strings from English to Somali. The input consisted of structured XML-like string elements used in Android applications, where each text label has a corresponding translation. ChatGPT provided translations for a comprehensive set of strings. The developer shared these translations in the pull request, leveraging ChatGPT’s ability to convert text into Somali while maintaining cultural and contextual relevance. However, the pull request was closed without merging with no explicit reason documented in the pull request discussion.
[…]
5.3.1 Implications for RQ4
Our findings indicate that pull request closures involving ChatGPT-assisted contributions are frequently influenced by project alignment, maintainability, and procedural requirements. These insights have important implications for both practitioners and researchers, particularly in understanding how AI influences pull request outcomes and what factors contribute to closure decisions.
[…]
While pull request closures occur in both AI-assisted and non-AI-assisted contributions, ChatGPT-generated patches introduce specific challenges due to their contextual misalignment and lack of repository awareness. To improve AI-assisted contributions and minimize pull request closures, developers should focus on the following key areas:
Assess maintainability and project alignment: Many AI-generated patches optimize performance, but do not fit long-term project goals. Developers should critically evaluate whether the suggestions of ChatGPT align with the architectural and maintainability requirements of the software before submission.
Ensure compliance with repository policies: pull requests were often closed due to missing procedural requirements, such as incorrect branch policies or lack of testing. Developers should manually verify that AI-generated patches conform to project guidelines.
Check for redundant contributions: AI-generated solutions sometimes duplicated existing efforts. Developers should review open pull requests and repository history before submitting an AI-assisted patch to prevent unnecessary duplication.
Engage with maintainers early: Some AI-assisted pull requests were closed without detailed feedback, indicating possible misalignment with the project goals. Developers should communicate with maintainers before submission to clarify expectations and increase the likelihood of acceptance.
[…]
Research should focus on addressing the specific challenges observed in AI-assisted pull request closures, particularly in areas where ChatGPT-generated contributions frequently fail. Key directions include:
Enhancing AI’s repository awareness: ChatGPT lacks knowledge of project-specific guidelines, leading to pull request rejections due to non-compliance with repository policies. Research should explore techniques for integrating AI with real-time repository metadata to improve adherence to project standards.
Improving AI-generated code maintainability: pull requests were frequently closed due to misalignment with the long-term project architecture. Future research should focus on adaptive AI models that analyze project history and architecture to generate more context-aware patches.
5.4 Implications for Educators
The findings of our study provide valuable insights for educators integrating generative AI tools such as ChatGPT into computer science curricula. By analyzing both merged and closed pull requests, along with varying patch integration levels, we highlight key learning opportunities. The classification into PA, PN, and NE cases offers distinct educational value: PA instances demonstrate AI-assisted coding in real-world, with varying integration percentages illustrating how developers selectively apply AI-generated code. PN cases encourage students to critically assess AI suggestions, understanding why patches may be rejected due to maintenance concerns or project-specific constraints. NE cases emphasize AI’s role beyond code generation, offering insights into conceptual learning, debugging strategies, and design discussions.
Additionally, our analysis of closed pull requests highlights the broader implications of AI-assisted contributions. Many pull requests were closed due to scope misalignment, maintainability concerns, or procedural issues, demonstrating that AI-generated solutions must align with long-term project goals and organizational constraints. Educators can use these examples to teach students the importance of human oversight, team collaboration, and reviewer feedback when working with AI-assisted code. This perspective helps students develop a more comprehensive understanding of software engineering workflows beyond patch integration.
These insights are actively applied in senior design projects at the University of Nevada Las Vegas (UNLV), where students analyze real-world instances of ChatGPT-assisted contributions. Case studies from this research are integrated into the curriculum, allowing students to engage in collaborative software development while exploring the ethical use of AI. ChatGPT is utilized for code review, debugging, refinement, and documentation, aligning with industry practices. Beyond senior design projects, these findings inform broader coursework: PN cases help students develop critical evaluation skills in programming courses, while NE cases illustrate how AI supports problem solving and iterative learning in software development. By studying both successful and rejected contributions, students gain a balanced perspective on the strengths and limitations of AI, preparing them for responsible AI-assisted development.
6 Related Work
Pull request decision-making: In collaborative software engineering and open-source projects, the decision-making process for pull requests is crucial to determine both the quality and the efficiency of the development. Gousios et al. [6] and Yu et al. [81] have examined how contributions are evaluated, highlighting the importance of peer review and social interactions. Various factors influence pull request acceptance, including technical merit, code quality, and alignment with project goals. For example, Tsay et al. [82] found that reviewers prioritize maintainability, readability, and adherence to coding guidelines when evaluating contributions. Similarly, Zhao et al. [20] showed that pull requests that lack comprehensive tests are more likely to be rejected, reinforcing the importance of thorough validation.
Beyond technical factors, social and ecosystem-wide dynamics also influence pull request decision-making. Dey and Mockus [22] examined how technical attributes and the reputation of contributors affect the acceptance of pull requests in the NPM ecosystem, highlighting the importance of project constraints and community participation. Storey et al. [14] further emphasized that communication dynamics across different channels profoundly shape participatory cultures in software development, underscoring the collaborative challenges that arise in distributed workflows such as pull requests. In addition to social factors, prior research has highlighted the value of mining developer disclosures to better understand development practices. Yu et al. [83] introduced Jitterbug, a machine learning-based approach for detecting self-admitted technical debt (SATD) in code comments. Inspired by this idea, our study extends the concept of self-admitted artifacts to AI-assisted development by analyzing self-admitted Chat
Meanwhile, Azeem et al. [21] proposed a classification-based recommendation system to predict pull request actions (accept, respond, or reject), helping integrators make more informed decisions within pull-based workflows. Our study builds on this body of research [84, 85, 86, 87] by extending the focus to AI-assisted pull request workflows, examining how ChatGPT influences pull request outcomes. By analyzing merged and closed pull requests, we provide empirical insights into how developers integrate, modify, or reject AI-generated patches, offering guidance on optimizing generative AI tools for improving pull request quality and accelerating decision-making.
AI-Assisted Software Engineering: Research on AI-assisted software engineering has examined the integration of ChatGPT-generated code in open-source projects, with studies such as Grewal et al. [33] and Siddiq et al. [34] providing key insights. Grewal et al. analyzed 3,044 ChatGPT-generated code snippets in GitHub repositories, finding that 54% of AI-generated code is retained in projects with minimal modifications. However, their study primarily measured retention rates and did not investigate the decision-making process behind the acceptance or rejection of the pull request. Similarly, Siddiq et al. evaluated the quality and security of ChatGPT-generated code, finding that AI-generated patches were rarely merged (only 12%) due to missing documentation, undefined variables, and security vulnerabilities such as improper resource management and hard-coded credentials. Recent work by Ehsani et al. [88] highlights that gaps in developer prompts can lead to ineffective LLM outputs, further complicating the adoption of AI-generated code in collaborative workflows. Preliminary work by Latendresse et al. [89] assessed ChatGPT’s effectiveness in recommending software libraries, raising concerns about licensing issues and integration reliability; however, their focus remains on library usage rather than collaborative coding workflows. Our study builds on these findings by analyzing why ChatGPT-generated patches are accepted, modified, or rejected within pull request workflows, providing a qualitative perspective on AI-assisted decision-making.
Beyond studies on AI-generated code, prior work has investigated how automation influences developer workflows and decision-making. Cassee et al. [90, 91] showed that automated systems such as Continuous Integration (CI) and GitHub bots alter code review behaviors and are critically evaluated by developers based on trust and perceived utility. Extending this line of inquiry, Shihab et al. [92, 93] analyzed the adoption of LLM-based chatbots and automation bots in pull request workflows, emphasizing challenges related to trust, accuracy, and context sensitivity. These findings underscore the need to examine how developers interact with suggested ChatGPT patches, critically assess them, and integrate or modify AI-assisted contributions during collaborative development.
In addition to studies on AI-generated code retention and quality, researchers have broadly examined AI-assisted software development, including the roles of ChatGPT and GitHub Copilot in code reviews, debugging, and software documentation [27, 28, 29, 94, 95, 96, 9]. Although these studies demonstrate improvements in productivity and code quality, they often focus on controlled experiments rather than real-world open-source development [30, 31, 32]. Previous work has explored the role of AI in pull request workflow automation [18, 35], but little research has examined how developers interact with ChatGPT in pull request discussions and adapt AI-generated patches to fit project-specific constraints. Our work addresses this gap by empirically analyzing ChatGPT’s impact in pull request workflows, particularly the decision-making processes that influence patch acceptance, modification, or rejection.
Prior research on software evolution and API maintenance has also shown that developers often face challenges when integrating changes due to harmful API usage patterns and breaking changes. Ochoa et al. [97] systematically characterized harmful API uses and available repair techniques, while Ochoa et al. [98] demonstrated how semantic versioning violations and API instability can complicate integration efforts. These challenges parallel the issues we observe when developers critically assess and adapt ChatGPT-suggested patches to fit evolving project architectures and avoid integration risks.
Code clone detection: PatchTrack uses clone detection techniques to determine if a patch was applied, aligning our work with prior studies on clone detection [99, 100, 50, 53, 101]. Baker [102] introduced a system for detecting verbatim code copying through text-based clone detection, while Kamiya et al. [103] developed CCFinder, a token-based tool that converts code into token sequences for similarity comparison—a method also employed in tools like Sourcercc [104], Rebedug [53], PaReco [50], and VUDDY [105]. Baxter et al. [106] demonstrated the use of abstract syntax trees (ASTs) for clone detection by identifying functional similarities despite syntax differences, as further emphasized by Lingxiao et al.’s DECKARD [107]. More recent studies leverage graph analysis and machine learning; for instance, Bowman et al.’s VGRAPH [108] uses code property triplets, while Feng et al. [109] combine token-based and ML techniques, reporting improved performance over traditional approaches. Beyond general-purpose clone detection, research on framework-specific clone analysis has provided insights into software ecosystems. Businge et al. [110] and Kawuma et al. [111] studied clones in the Eclipse framework, reporting that clone occurrences range between 9% and 10%. These findings illustrate the prevalence of duplicated code even in structured development environments, reinforcing the importance of precise clone detection techniques, such as those implemented in PatchTrack, to enhance software quality and maintainability.
7 Threats to Validity and Trustworthiness
7.1 Threats to Validity (RQ1)
For RQ1, the primary threat to validity lies in the accuracy of PatchTrack. Reliance on n-gram tokenization with can introduce false positives, and normalization steps could inadvertently alter the structure of some patches. Furthermore, PatchTrack does not capture structural or semantic differences between ChatGPT-generated patches and applied code, as it focuses on exact token matching rather than deeper syntactic or semantic equivalences. Although this approach is effective for detecting exact matches, it can classify substantially modified patches as PN (patch not applied), even when ChatGPT influenced the final implementation. To mitigate this, we manually validated a subset of cases and refined our classification approach, though some edge cases remain.
Another limitation stems from the representativeness of the dataset. Our analysis is restricted to pull requests where developers explicitly shared ChatGPT-generated content, meaning it does not capture private AI interactions or cases where ChatGPT was used without documentation in pull request discussions. This could introduce selection bias, as our findings are based only on visible ChatGPT contributions.
7.2 Trustworthiness of Qualitative Findings (RQ2, RQ3, RQ4)
For RQ2, RQ3, and RQ4, we assess the trustworthiness of our qualitative findings using Lincoln and Guba’s framework [112].
Credibility
To ensure credibility, we applied the Framework Method and thematic extraction via card sorting. Manual classification was performed in 277 cases ( PA, PN, NE, and CL), with Cohen’s Kappa of 0.85, indicating strong agreement among reviewers. Verbatim quotes from pull request discussions support data-driven interpretations, reducing researcher bias.
Transferability
Although our findings are derived from GitHub pull request discussions, they may be applicable to other AI-assisted software development workflows. We provide detailed descriptions of our dataset, coding process, and themes, allowing future researchers to assess the applicability of our findings to different contexts.
Dependability
We document all coding procedures, classification criteria, and iterative discussions to refine the definitions of the themes. This ensures consistency in analysis and allows replication in future studies.
Confirmability
To minimize researcher bias, thematic extractions followed structured decision rules and were cross-verified using independent coding rounds. Direct excerpts from pull request conversations ensure that interpretations are grounded in actual data.
7.3 Limitations and Future Work
One limitation of our study is that we did not differentiate between the versions of ChatGPT used in the analyzed conversations. As newer versions improve accuracy and reasoning, different versions may have influenced developer engagement and patch integration. However, our dataset does not capture explicit versioning information, making it difficult to assess these differences. Future research could explore the impact of ChatGPT versioning on patch acceptance and developer adaptation strategies.
Additionally, while our dataset spans multiple projects, we did not analyze project-level differences such as repository activity or programming language. Future work could examine whether these factors influence ChatGPT-assisted patch integration.
Another limitation comes from the representativeness of the dataset. Our analysis is restricted to pull requests where developers explicitly shared ChatGPT-generated content, which we classify as self-admitted ChatGPT usage (SACU). This means that it does not capture private AI interactions or cases where ChatGPT was used without documentation in pull request discussions. This could introduce selection bias, as our findings are based only on visible ChatGPT contributions.
Beyond these limitations, future research should explore ways to improve AI-generated recommendations by incorporating project-specific constraints, enhancing maintainability alignment, and refining AI-assisted workflows. Investigating the long-term impact of ChatGPT-assisted patches on software quality and maintainability could also provide insights into whether frequent AI usage leads to sustained code improvements or introduces new challenges related to technical debt.
8 Ethical Implications
Ethical integrity is paramount in research utilizing public platforms like GitHub. While repositories are publicly accessible, privacy considerations remain crucial as our analysis involves examining pull requests without explicit consent. To address this, we ensured proper attribution by explicitly acknowledging all pull requests that contributed to our study and discussion. This includes citing relevant pull requests in our references and, where applicable, linking directly to their commits to appropriately credit the contributors.
To promote transparency and reproducibility, we have made our replication package publicly available, including datasets, analysis scripts, and supplementary documentation[39]. These measures uphold the ethical responsibilities of using open-source data while fostering reproducibility and openness in computer science research.
9 Conclusion
This study examined ChatGPT’s role in patch decision-making within pull requests by analyzing 285 merged pull requests and 53 closed pull requests from 255 repositories. We classified ChatGPT-generated patches as applied (PA), not applied (PN), or not suggested (NE), identifying 115 PA, 64 PN, and 106 NE cases. The median integration rate of 25% suggests that AI-generated patches are often treated as starting points rather than final implementations, and developers selectively modify or discard parts to fit project needs. Analysis of PN cases revealed six themes in which developers adapted AI suggestions to fit project constraints, while cases NE highlighted ChatGPT’s broader role in debugging, documentation, and conceptual guidance.
Our findings indicate that while ChatGPT accelerates software development by providing structured suggestions, human oversight remains critical in refining AI-generated contributions. Closed pull requests showed that rejections were often due to scope misalignment, maintainability concerns, or procedural barriers, underscoring the challenges of integrating AI into collaborative workflows. Overall, this study provides empirical insights into how generative AI is used in pull request workflows, informing developers, maintainers, and educators on best practices for AI-assisted coding and decision-making.
Declarations
Funding
This research was supported by the NEVADA NASA EPSCoR program under Award No. AWD2536/GR20392.
Ethics approval
Not applicable
Informed consent
Not applicable
Author contribution
John Businge: Conceptualization, supervision, and review and editing of the manuscript.
Daniel Ogenrwot: Data collection, analysis, and writing of the original draft.
Data availability
The datasets, analysis scripts, and supplementary documentation are publicly available at https://zenodo.org/records/14978625
Conflict of interest
The authors declare that there are no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Clinical trial number in the manuscript
Not applicable
Appendix A Supplementary Tables
| Classification Type | n = 1 | n = 2 | n = 3 | n = 4 |
| Patch Applied (PA) | 116 | 83 | 47 | 38 |
| Patch Not Applied (PN) | 63 | 96 | 132 | 141 |
| No Patch Exists (NE) | 106 | 106 | 106 | 106 |
| N-Gram | n = 1 | n = 2 | n = 3 | n = 4 | ||||||||||||
| Acc. | Prec. | Recall | F1 | Acc. | Prec. | Recall | F1 | Acc. | Prec. | Recall | F1 | Acc. | Prec. | Recall | F1 | |
| PA | 97.2 | 96.6 | 95.6 | 96.1 | 95.6 | 95.3 | 88.4 | 91.7 | 95.9 | 88.1 | 88.1 | 88.1 | 96.7 | 90.9 | 85.7 | 88.2 |
| PN | 95.4 | 92.9 | 94.5 | 93.7 | 95.3 | 90.3 | 96.2 | 93.2 | 95.9 | 94.9 | 94.9 | 94.9 | 94.5 | 95.2 | 97.5 | 96.3 |
| NE | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 |
| All | 97.5 | 96.5 | 96.7 | 96.6 | 96.9 | 95.2 | 94.7 | 94.9 | 97.2 | 94.3 | 94.3 | 94.3 | 97.1 | 95.3 | 94.4 | 94.8 |